A U.S. company has a ¥750 million payable due in one year to a bank in Japan. The current spot rate S($/¥) = $0.0086/ ¥ and the one - year forward rate F 360 ($/¥)= $0.0092/¥. The annual interest rate is 3 percent in Japan and 6 percent in the United States.

a. How to implement a hedge using a forward contract? Compute the guaranteed dollar payment in one year using the forward hedge.

b. How to implement a money market hedge? Compute the guaranteed dollar payment in one year using money market hedge.

c. If the U.S company decides to hedge using call option on yen, what would be the dollar payment in one year? Assume that the spot exchange rate S($/¥) = $0.0092/¥ in one year. Also suppose that a one-year call option on yen has an exercise price of $0.0086/¥ and a premium of $0.00012/¥.
Answer #1

PT a Yen is payable in 1 year so the US company has risk of Yen rising. In order to hedge , US company need to buy forward contract.

Dollar payment = Yen 750 million*$0.0092 per Yen =$ 6.9 million payable

PT b

To implement money market hedge we need to borrow amount In USD and then convert it into Yen and then this amount is Invested in Yen for 1 year.

Yen Payable =750 million

Present value of Yen = 750/(1.03) i.e. Yen 728.15 million

Dollar required for Buying Yen 728.15 million =Yen 728.15 * 0.0086 i.e.USD 6.26209million

Borrow amount in USD 6.26209 million , Interest =6.26209 million *6% i.e. USD 0.37572 million

Total Dollar payable =6.26209 million +0.37572 million i.e. USD 6.63781 million

PT c Expected rate =0.0092/Yen

Strike price = 0.0086/Yen

  Premium for call = 0.00012/Yen

Effective cost = Strike price + Premium

= 0.0086+0.00012 i.e. 0.00872/Yen

Dollar payable =Yen 750 million *0.00872 i.e. USD 6.54 million